Directions: Analyze the following equations. Then answer the given questions.

1 4FeS + 70 2 2Fe 2 O 3 + 4SO 2 Calculate the number of moles of Fes if there are 10 moles of Fe,O, 2 BCO + 17H 2 C 8 H 18 + 8H 2 O Calculate the number of moles of C_{1}*H_{10} there are 40 moles H_{2} ?

3. 2As + 6NaOH 2Na,AsO 3 + 3H 2 Calculate the mole of As if there are 6 moles of NaOH?

4. TiCl 4 + 2H 2 O TiO 2 + 4HCl Calculate the mole of if there are 4 moles of H_{2}*O ?

5. 4NH 3 + 5O 2 4NO + 6H 2 O Calculate the mass of NO if there are 4 moles of N*H_{3} ,?


1
Expert's answer
2021-10-28T02:57:57-0400

1. 4FeS + 7O2 → 4SO2 + 2Fe2O3

n(FeS) = 2n(Fe2O3) "= 2 \times 10 = 20 \;moles"

2. 8CO + 17H2 → C8H18 + 8H2O

n(C8H18) = "\frac{1}{17}" n(H2) "= \frac{1}{17} \times 40 = 2.35 \;mol"

3. 2As + 6NaOH → 2Na3AsO3 + 3H2

n(As) = "\frac{1}{3}" n(NaOH) "= \frac{1}{3} \times 6 = 2 \;mol"

4. TiCl4 + 2H2O → 4HCl + TiO2

n(TiCl4) = "\frac{1}{2}" n(H2O) "= \frac{1}{2} \times 4 = 2 \;mol"

5. 4NH3 + 5O2 → 4NO +6H2O

n(NO) = n(NH3) = 4 moles

MM(NO) = 30 g/mol

m(NO) "= 4 \times 30 = 120 \;g"